how can i solve this problem using the elimination method for systems x-6y=44 2x-7y=53
Answer by mananth(16946)

x-6y=44 .............1
2x-7y=53.............2
Eliminate y
multiply (1)by 7
Multiply (2) by -6
7x-42y= 308
-12x+42 y=-318

Add the two equations

-5x=-10

/-5

x=2

plug value of x in (1)

x-6y=44
2-6y=44
-6y=44-2
-6y=42
y=-7